OpenStudy (anonymous):

A pair of dice is rolled. What is the probability that the sum of the numbers rolled is either an even number or a multiple of 3?

OpenStudy (anonymous):

I have the answer, and I mostly know how to do it; I just have one question.

OpenStudy (anonymous):

Because it is not mutually exclusive, to calculate this I use this P(even) + P(multiple of 3) - P(6) But according to the site I got this there are 6 ways of rolling a six, not 5 which is what I got

OpenStudy (anonymous):

6 1-5 2-4 3-3 4-2 5-1

OpenStudy (anonymous):

Unless it is counting 3-3 twice

OpenStudy (anonymous):

Am I supposed to count 3-3 twice?

OpenStudy (mathstudent55):

I don't think so. there is only one way to roll a 3-3

OpenStudy (anonymous):

This is their explaination: Of the 36 possible outcomes, 18 are even sums. P(even) = 18/36 = 1/2 Sums of 3, 6, 9, and 12 are multiples of 3. There are 12 sums that are multiples of 3. P(multiple of 3)= 12/36 = 1/3 However, some of these outcomes appear in both events. (not mutually exclusive) The sums that are even and a multiple of 3 are 6 and 12. There are 6 ordered pairs with these sums. P(even AND a multiple of 3) = 6/36 = 1/6 P(even OR a multiple of 3) = 18/36 + 12/36 - 6/36 = 24/36 = 2/3

OpenStudy (anonymous):

Oh sorry

OpenStudy (anonymous):

Probability of rolling a 6 AND 12, forgot about the twelve

OpenStudy (anonymous):

Whats the difference between mutually inclusive and mutually exclusive?

OpenStudy (mathstudent55):

Mutually exclusive means both events can't happen at the same time.

OpenStudy (anonymous):

Thank you, I think I understand the difference.

OpenStudy (mathstudent55):

A silly eample of mutually exclusive is if you were asked to pick a number from 1 to 10. What's the probability it will be an odd number of a multiple of 2. There are 5 odd numbers from 1 to 10 and there are 5 multiples of 2 from 1 to 10. You add the probabilities The two choices are mutually exclusive bec if a nunmber is a multiple of 2, it is even, and it can't be odd.

OpenStudy (mathstudent55):

You're welcome.